OpenStudy (ghostedly):

Hey guys... this is SO confusing... What does n stand for in 2(-7)/2^n=2^2

OpenStudy (anonymous):

its the question

OpenStudy (ghostedly):

Here, http://prntscr.com/8r1k22

Nnesha (nnesha):

\[\frac{ 2^{-7} }{ 2^n}=2^2\]is this ur question ?

Nnesha (nnesha):

oh nvm

OpenStudy (anonymous):

\[\frac{ -14 }{ 2^{n} }= 2^{2}\]

Nnesha (nnesha):

solve for n ?? hmm

Nnesha (nnesha):

hmm what's the topic ?? log or ln ??

OpenStudy (jdoe0001):

ditto

OpenStudy (jdoe0001):

what chapter is this covering? or chapter section

Nnesha (nnesha):

bitto

OpenStudy (anonymous):

and, he gone

OpenStudy (anonymous):

\[2^n=\frac{ -7 }{ 2 }\] i think no real value of n will give this result.

OpenStudy (jdoe0001):

hehe, guess it was just a dry-by

OpenStudy (ghostedly):

Hm